OpenStudy (anonymous):

sqrt 4a+1 endsqrt +3=2

OpenStudy (anonymous):

\[\sqrt{4a+1}+3=2\]? if so no solution

OpenStudy (anonymous):

if you start by subtracting \(3\) you get \[\sqrt{4a+1}=-1\] and the radical on the left is greater than or equal to zero, so it cannot be \(-1\)

OpenStudy (anonymous):

so its a null set correct?